### 09 Apr 2026 v1

# Top Management Journal Portal: A Real-Source Search and Research Analytics Artifact for UTD-24 and FT50 Journals

Abstract
This paper presents Top Management Journal Portal, a deployable web artifact for searching, monitoring, and interpreting literature from elite business and management journals. The system integrates the UTD-24 and Financial Times 50 (FT50) journal pools, retrieves live article metadata from the Crossref REST API, and organizes scholarly work into an end-to-end workflow spanning query formulation, result filtering, hotspot extraction, citation export, favorites management, and usage analytics. Unlike static journal directories or general-purpose academic search engines, the artifact is explicitly scoped to high-status management outlets and is designed to support sensemaking tasks that matter to researchers, doctoral students, and lab managers: identifying recent work, surfacing topical concentration, and converting search output into actionable research material. Architecturally, the system emphasizes source transparency, modularity, and low-cost public deployability through a lightweight Node.js service layer, a multi-page client interface, optional large-language-model enhancement for hotspot rewriting, and a free-tier persistence path through Supabase. The paper contributes both a functioning design artifact and an extensible architectural pattern for journal-pool-specific scholarly discovery, with implications for digital research infrastructure in information systems and business scholarship.
Keywords: design science, scholarly search, research analytics, business journals, Crossref, digital research infrastructure
1 Introduction
Researchers in business schools often rely on elite journal lists as practical filters for literature review, outlet targeting, and academic evaluation. Two such lists are especially influential: the UTD-24, commonly used in research productivity assessment, and the FT50, frequently referenced in ranking, accreditation, and prestige signaling contexts. Yet the working process around these lists remains fragmented. Journal pools are typically maintained separately from live literature search, publisher pages provide inconsistent browsing experiences, and monitoring topical developments across multiple elite journals is labor intensive.
This paper introduces Top Management Journal Portal, a web-based research artifact designed to reduce that fragmentation. The artifact does not attempt to replace publisher websites or commercial bibliographic databases. Rather, it provides a focused, domain-specific decision support layer that combines journal-pool curation, real-source metadata retrieval, topic extraction, export utilities, and lightweight behavioral analytics. In design-science terms, the artifact addresses a class of recurrent scholarly search and sensemaking problems by creating a practical digital object for use in a specific problem domain (Hevner et al., 2004; Gregor and Hevner, 2013). The broader ambition is to show that even a relatively lightweight artifact can offer meaningful infrastructure value when its scope, sourcing logic, and task support are carefully aligned.
The current public instance of the artifact is available online [^1] https://utd24-journal-portal.onrender.com/, which allows the system to function not only as a conceptual design exemplar but also as a live, inspectable research artifact.
The artifact was designed around three overarching objectives. First, it should help researchers quickly search within a trusted set of top management and business journals. Second, it should provide interpretable signals about emerging topics from returned results rather than merely displaying lists of papers. Third, it should remain easy to deploy and maintain in low-budget, lab-managed, or publicly hosted environments.
The paper proceeds as follows. Section 2 positions the artifact in design science and scholarly discovery work. Section 3 outlines design objectives and artifact requirements. Section 5 presents the architecture and search-to-insight workflow. Section 6 describes the research analytics layer. Section 7 discusses implementation and deployment decisions. The final sections discuss contributions, limitations, and future directions.
2 Background and Research Motivation
Design science research in information systems emphasizes the creation and evaluation of artifacts that address relevant organizational and human problems (Hevner et al., 2004). In this tradition, the artifact presented here is best understood as a domain-specific search and sensemaking system for scholarly work. It occupies the intersection of information retrieval, research support, and digital infrastructure for academic decision making. The design also resonates with the view that information systems artifacts become more valuable when they embed both operational functionality and interpretive support (March and Smith, 1995; Gregor and Hevner, 2013).
The motivating problem is not a lack of journal information per se. Instead, the problem lies in the separation between (1) curated notions of journal quality and prestige, (2) real-time metadata about recent papers, and (3) tools that help scholars quickly interpret what a search result set means. General search engines offer breadth, but they are not optimized for journal-pool-specific exploration. Static journal directories offer curation, but rarely offer live query-based insight generation. This artifact addresses that gap through a lightweight but integrated workflow. From a research-infrastructure perspective, the artifact is promising because it reduces the translation cost between “finding papers” and “understanding what a query reveals about a field.”
This positioning is also consistent with prior work that compares scholarly search and citation infrastructures across coverage, retrieval quality, and analytical usefulness. Comparative studies have shown that Google Scholar, Web of Science, Scopus, PubMed, and other academic search systems differ materially in coverage, precision, transparency, and suitability for specific scholarly tasks (Bakkalbasi et al., 2006; Falagas et al., 2008; Harzing and Alakangas, 2016; Martín-Martín et al., 2018; Gusenbauer, 2022). In parallel, scientometric work has demonstrated the value of visual and structural techniques for detecting emerging topics and research fronts in large document collections (Chen, 2006). The present artifact contributes to this broader conversation by operationalizing a journal-pool-specific layer for management research rather than attempting universal scholarly search.
Table 1: Problem Context and Artifact Response
Problem Area Observed Limitation Artifact Response
Elite journal navigation UTD-24 and FT50 are consulted separately; outlet information is fragmented Unified journal-pool directory with pool membership, official links, and submission links
Current literature discovery Researcher must scan multiple journal or publisher pages manually Live query-based retrieval from Crossref constrained by journal pools
Topic sensing Search results alone do not reveal topical concentration or method signals Hotspot extraction, distributions, and result-set summaries
Practical research workflow Search, export, citation formatting, and bookmarking are disconnected Citation modal, favorites, DOI links, and CSV/report export
Public deployment and maintenance Low-cost hosting often breaks persistence and increases maintenance burden Minimal Node.js stack with optional Supabase-backed persistence
3 Design Objectives
The artifact was built around five design objectives.
Domain focus. The portal is intentionally limited to journals in the UTD-24 and FT50 pools. This restriction reduces noise and aligns the interface with business-school research evaluation logic.
Real-source transparency. All journal and article content is tied to identifiable sources. Journal pool membership is documented, article metadata are retrieved from Crossref, and output records expose DOI and source links.
Query-to-insight workflow. The system should not only retrieve search results but also help users interpret them through hotspot analysis, distributions, affiliation views, and simple trend analytics.
Practical research support. Citation export, favorites, downloadable CSV output, and quick article-level commentary are included because they reduce friction in everyday research tasks.
Low-cost deployability. The artifact should remain easy to deploy without a heavy dependency chain, while still supporting persistent analytics on a free hosting path.
Table 2: Core Design Objectives and Implemented Features
Design Objective Operationalization Implemented Features
Domain focus Restrict scope to elite business journals UTD-24/FT50 pools, list filters, journal directory
Source transparency Expose provenance at the object level DOI links, source notes, official site links, submission links
Search-to-insight Transform query results into interpretable summaries Hotspots, distributions, affiliation ranking, method signals
Workflow support Reduce post-search friction Citation modal, BibTeX export, favorites, CSV export
Deployability Keep infrastructure lean and maintainable Node.js server, static front end, Render deployment, Supabase persistence
4 Artifact Overview
The artifact follows a lightweight client-server architecture. The front end is a modular multi-page website; the back end is a small Node.js application that serves static assets, proxies metadata retrieval, computes analytics, and persists usage events. The current implementation is intended as a publicly deployable research artifact rather than a large-scale platform. This design choice is deliberate: the goal is to demonstrate that a comparatively small technical stack can still support a meaningful research workflow when the architecture is organized around clear scholarly tasks and traceable data provenance.
4.1 Front-End Structure
The user-facing interface is organized into five modules:
  1. 1.
    Home: overview, source note, and quick-access entry points.
  2. 2.
    Journal Directory: filtered browsing of UTD-24 and FT50 journals.
  3. 3.
    Article Search: Crossref retrieval, filters, hotspot analysis, export, favorites, and citation support.
  4. 4.
    Resources: conferences, submission tools, and external scholarly resources.
  5. 5.
    Analytics: visits, top search keywords, favorites ranking, trends, and traffic map widgets.
4.2 Back-End Services
The Node.js back end performs four roles:
  1. 1.
    static file serving,
  2. 2.
    metadata retrieval and normalization,
  3. 3.
    analytical summary generation,
  4. 4.
    usage analytics storage and reporting.
Table 3: Front-End and Back-End Technical Composition
Layer Implemented Technology Technical Role in the Artifact
Presentation layer HTML, CSS, vanilla JavaScript Multi-page interface for search, journal browsing, export utilities, and analytics views
Search interaction Browser-side form state and query serialization Stable query submission, shareable URLs, local favorites, and citation actions
Visualization layer ECharts Hotspot ranking, distributional views, and cumulative traffic analytics
Application layer Node.js HTTP server Static asset delivery, API routing, Crossref orchestration, and result normalization
Metadata source Crossref REST API Live retrieval of paper-level metadata, dates, authors, affiliations, and DOI records
Persistence layer Supabase or local JSON fallback Analytics durability across deployments and lightweight operational monitoring
Deployment layer Render-compatible hosting path Public accessibility with minimal infrastructure overhead
Presentation Layer Home    Journal Directory    Article Search    Resources    Analytics Application Layer Node.js server for routing, Crossref request orchestration, metadata normalization, query filtering, pagination, citation formatting, and analytics aggregation Analytics and Enrichment Layer Hotspot extraction    distributional analytics    citation export    favorites ranking    optional LLM enhancement Curated Sources UTD-24 metadata FT50 metadata External Metadata Crossref REST API Persistence Supabase analytics store or local fallback
Figure 1: System architecture of the Top Management Journal Portal artifact
5 Search-to-Insight Workflow
The portal implements a query-to-insight workflow rather than a search-only pipeline. Figure 2 summarizes the operational sequence.
Select Journal PoolUTD-24 / FT50 / AllEnter Queryand FiltersIssue BoundedCrossref RequestsNormalize MetadataTitle, DOI, Authors, DateFilter, Sort,and PaginateCompute Hotspotsand AnalyticsRender Articles,Charts, and Exports
Figure 2: Search-to-insight workflow
Given a keyword and optional filters, the back end first identifies the journals that belong to the selected pool. It then issues bounded Crossref requests journal by journal, normalizes the returned data into a common article schema, applies filtering and sorting logic, and computes analytical summaries from the resulting set.
This workflow is important because it turns journal-pool curation into an operational search boundary. Rather than querying a global metadata space and later attempting to infer whether results belong to top journals, the system begins from a trusted elite-journal pool and then builds the result set upward.
The implemented interface focuses on a small number of stable research tasks: choosing a journal pool, launching a live query, inspecting recent papers, extracting hotspots, and exporting references. Instead of emphasizing low-fidelity mockups, the report highlights the implemented views that researchers actually interact with during search and analysis. Figures 35 therefore serve not only as screenshots, but as evidence of how the artifact translates architectural intentions into concrete researcher-facing workflows.
Figure 3: Implemented homepage view showing real-source positioning, quick access modules, and statistics
Figure 4: Implemented search interface showing live article retrieval, hotspot analysis, citation tools, and filtering controls
Figure 5: Implemented analytics interface with trend views, source tracking, favorites ranking, and traffic visualization
Figure 3 illustrates the artifact’s front-door logic: it foregrounds source transparency, journal-pool framing, and navigational shortcuts. Figure 4 shows the core operational surface, where live metadata retrieval, filtering, citation export, and hotspot analysis are tightly coupled in a single page. Figure 5 extends the artifact beyond search by exposing a monitoring layer for visits, usage concentration, and interaction patterns. Together, these implemented views demonstrate that the artifact is not only technically functional but also structurally aligned with recurring scholarly work practices.
6 Research Analytics Layer
The research analytics layer is the artifact’s most distinctive feature. It aims to support rapid scholarly orientation rather than formal bibliometric inference. In other words, the portal is designed less as a replacement for specialized bibliometric software and more as a decision-support layer that helps researchers recognize what is salient in a query result set quickly enough to guide the next step of reading, downloading, or theorizing.
6.1 Hotspot Extraction
Hotspot phrases are derived from titles and abstracts of the current result set. Candidate phrases are scored and then grouped into interpretable categories such as AI/Algorithms, Consumer/Market, Operations/Supply Chain, Finance/Accounting, and Strategy/Organization. When no native article keywords are available, the system falls back to rule-based extraction. When an LLM endpoint is configured, hotspot labels and quick takes can be rewritten into more readable summaries.
6.2 Distributional Analytics
The portal computes a lightweight analytical layer from the current search result set:
  • journal distribution,
  • category distribution,
  • year distribution,
  • keyword distribution,
  • top affiliations,
  • method signals,
  • top cited papers,
  • abstract and affiliation coverage.
These summaries provide immediate situational awareness for users exploring unfamiliar topics across multiple elite journals. This matters because management researchers often work across fragmented topical boundaries; a lightweight but interpretable analytic layer can materially reduce the cognitive effort required to move from article retrieval to exploratory synthesis.
D-9D-8D-7D-6D-5D-4D-3D-2D-1D00101020203030Recent 10-Day WindowCumulative VisitsExample cumulative traffic trend
Figure 6: Illustrative cumulative analytics view used in the portal dashboard
6.3 Behavioral Analytics
The artifact also records page visits, search keywords, and favorites events. These events are aggregated into top pages, source distributions, most frequent search terms, and most favorited papers. For public deployment, a ClustrMaps widget can be attached to the analytics page to visualize geographic traffic sources. Although these measures are operational rather than theoretical, they create a useful substrate for future evaluation studies by revealing how scholars actually interact with the artifact over time.
Table 4: Analytics Objects and Intended Use
Analytics Object Computation Basis Primary Use
Hotspot phrases Title and abstract phrase extraction with scoring Rapid topic sensing within a query result set
Keyword distribution Crossref subject fields or fallback extraction Quick topical orientation
Top affiliations Aggregated author affiliation strings Institutional awareness and field mapping
Method signals Heuristic text pattern detection Rough understanding of empirical orientation
Top favorited papers Aggregated user bookmark events Monitoring researcher interest and repeated use
Visit trend and source mix Pageview events and referrer classification Artifact monitoring and public usage awareness
7 Technical Architecture and Implementation
7.1 Technology Stack
The artifact uses Node.js for server-side routing and API orchestration, vanilla JavaScript for front-end logic, HTML/CSS for multi-page rendering, ECharts for result-set and analytics visualization, Supabase for free-tier persistence of usage analytics, and Render as a practical public deployment target. The implementation deliberately avoids a heavier front-end framework in order to minimize setup friction and hosting complexity. This restraint is important for the artifact’s design logic: a research support system is more likely to be adopted, replicated, and extended when the technical barrier to deployment remains low.
7.2 Front-End and Interaction Design
The front end is organized as a modular multi-page interface rather than a monolithic single-page application. This choice simplifies routing, page ownership, and public hosting while preserving a coherent user journey across home, journal browsing, live search, resource access, and analytics. Browser-side logic manages form state, URL synchronization, favorites persistence, citation actions, and optional user-provided large-language-model configuration. From a software-engineering perspective, this structure improves maintainability by keeping task-specific rendering concerns loosely coupled while still allowing shared interface patterns across pages.
7.3 Back-End Retrieval and Normalization
The back end mediates between curated journal pools and live metadata retrieval. For each query, the service first resolves the relevant journal set, then issues bounded Crossref requests on a journal-by-journal basis, normalizes returned records into a common schema, removes duplicates, applies filters, and computes analytical summaries. This orchestration layer is central to the artifact’s contribution: it operationalizes prestige-based journal curation as a search boundary, thereby aligning retrieval behavior with the evaluative logic commonly used in business schools.
7.4 Data Provenance and Source Integrity
The artifact is intentionally conservative about data provenance. Journal pools are anchored in documented UTD-24 and FT50 sources, while paper-level content is derived from Crossref records and exposed through DOI links and source labels. When article keywords are unavailable from native metadata, the system falls back to explicit heuristic extraction rather than presenting opaque or untraceable enrichment. Optional LLM support is treated as a presentation enhancement layer rather than a source of authoritative bibliographic truth.
7.5 Persistence and Deployment
One implementation challenge concerns analytics persistence on low-cost hosting. Some free web-service environments provide ephemeral file systems, which would reset local counters on redeploy. The artifact therefore supports two persistence modes:
  1. 1.
    local JSON storage for local development and quick tests,
  2. 2.
    Supabase-backed analytics persistence for stable public deployment.
This decision is not only technical but also methodological. If the artifact is to be used as public research infrastructure, persistence reliability becomes part of the artifact’s practical validity. A deployable research artifact should not merely demonstrate conceptual fit; it should also remain stable enough for repeated public use, classroom demonstration, or lab-level extension.
7.6 Citation and Export Support
Each article object can be exported or cited through several output modes, including BibTeX, APA, MLA, Chicago, and plain text. CSV export and report export are also supported. These functions position the artifact closer to an everyday research tool than a demonstration-only prototype. In practice, they reduce a common break in scholarly workflow: the transition from discovering relevant papers to incorporating them into notes, bibliographies, and downstream review documents.
8 Discussion
The artifact contributes to design science and scholarly infrastructure in three ways. First, it demonstrates a domain-specific scholarly search artifact centered on elite business journals rather than undifferentiated bibliographic breadth. Second, it shows how real-source metadata and lightweight analytics can be combined into a usable search-to-insight workflow without requiring proprietary infrastructure. Third, it offers a deployable pattern for small research labs that need a manageable public-facing tool. More broadly, the artifact suggests that there is substantial room for middle-layer scholarly systems: systems that do not compete with large bibliographic platforms on scale, but outperform them on domain alignment, transparency, and task focus.
At the same time, the artifact has important limitations. Coverage depends on Crossref metadata availability, which varies across journals and articles. The hotspot layer is optimized for interpretability rather than formal topic-model validity. The system does not yet support collaborative annotation, full-text indexing, or citation-network exploration. Finally, the artifact has not yet undergone formal user evaluation with faculty or doctoral students. These limits should be interpreted as research opportunities rather than as disqualifying weaknesses: they point to the next design cycle needed to move from promising artifact to evaluated scholarly infrastructure.
9 Implications and Future Work
Several extensions are especially promising:
  • User evaluation: task-based studies with doctoral students and faculty to assess search efficiency, interpretability, and perceived usefulness.
  • Richer analytics: topic modeling, co-author networks, institution collaboration views, and cross-pool comparison logic.
  • Temporal intelligence: rolling topic evolution across years, pools, and journals, enabling dynamic field monitoring.
  • Submission support: connecting observed topic concentration to outlet targeting guidance and journal-fit recommendations.
  • Artifact generalization: adapting the same architecture to adjacent domains such as finance, information systems, operations, or interdisciplinary journal pools.
Taken together, these directions suggest that the portal can evolve from a useful search utility into a broader research-infrastructure artifact: one that supports discovery, interpretation, and eventually strategic research planning.
10 Conclusion
Top Management Journal Portal is a focused research artifact for discovering, searching, and analyzing literature from elite business journals. By integrating UTD-24 and FT50 curation with live Crossref retrieval, lightweight analytical summaries, optional AI enhancement, citation utilities, and deployable analytics infrastructure, the artifact offers a practical response to the fragmented search workflows often faced by management researchers. Although it remains an early-stage artifact rather than a finished scholarly platform, it already demonstrates a credible and extensible design pattern for business-school-oriented digital research infrastructure. For ICIS-style design and artifact discussions, the promise of the system lies not only in what it currently does, but in the evaluative and infrastructural research program it makes possible.
Acknowledgments
The current public-facing artifact credits Designed by Chuang Zhao (ADM Lab). The implementation also integrates external data and infrastructure services, including Crossref, Supabase, Render, and ClustrMaps.